# Robust Containment Queries over Collections of Trimmed NURBS Surfaces via Generalized Winding Numbers

### v2

@@ Appendix C Supplemental Material
In this supplemental material, we evaluate our 3D GWN algorithm qualitatively and quantitatively on several additional watertight and non-watertight CAD models. While most of these additional shapes are taken directly from the ABC dataset (Koch et al., 2019), others were either constructed by hand in modeling software or sourced from other online repositories.
For each shape, we provide summary statistics in the format of Table 1 from the primary manuscript: We compute the GWN over a uniform grid of $50\times 50\times 50$ points within the shown bounding box and report the average time, in milliseconds, needed to evaluate the GWN for each surface in the shape (overall and by case), alongside the breakdown of all point-surface GWN calculations by case. We also visualize the 3D GWN scalar field along different 2D slices, and show the nearest integer value to each scalar. These integers can be straightforwardly mapped to a containment decision using either a non-zero rule, or an even-odd rule.

### v3

@@ Appendix C Supplemental Material
In this supplemental material, we evaluate our 3D GWN algorithm qualitatively and quantitatively on several additional watertight and non-watertight CAD models. While most of these additional shapes are taken directly from the ABC dataset (Koch et al., 2019), others were either constructed by hand in modeling software or sourced from other online repositories. The specific models used in this work can be found in (Axom_data, 2025).
For each shape, we provide summary statistics in the format of Table 1 from the primary manuscript: We compute the GWN over a uniform grid of $50\times 50\times 50$ points within the shown bounding box and report the average time, in milliseconds, needed to evaluate the GWN for each surface in the shape, alongside the breakdown of all point-surface GWN calculations by case. We also visualize the 3D GWN scalar field along different 2D slices, and show the nearest integer value to each scalar. These integers can be straightforwardly mapped to a containment decision using either a non-zero rule, or an even-odd rule.